OpenStudy (calculator):

The equation of a curve is y = √(8x − x2). Find (i) an expression for dy dx , and the coordinates of the stationary point on the curve,

OpenStudy (calculator):

i get x=0 ,x=-4 , x=8, x=4

OpenStudy (kropot72):

Did you get the following for dy/dx? \[dy/dx=1/2(8x-x ^{2})^{-1/2}\times (8-2x)\]

OpenStudy (calculator):

\[y=(\sqrt{8x-x^2})\] \[y=(8x-x^2)^{\frac{1}{2}}\] \[y=\frac{1}{2}(8x-x^2)^{-1/2}(8-2x)\] \[4(8x-x^2)^{-1/2}-x(8x-x^2)^{-1/2}=0\] \[4(8x-x^2)^{-1/2}=x(8x-x^2)^{-1/2}\] \[\frac{4}{\sqrt{8x-x^2}}=\frac{x}{\sqrt{8x-x^2}}\] \[\frac{16}{8x-x^2}=\frac{x^2}{8x-x^2}\] \[128x-16x^2=8x^3-x^4\] \[x^4-8x^3-16x^2+128x=0\] \[x(x^3-8x^2-16x+128)=0\] x=0, x=-4,x=8,x=4

OpenStudy (calculator):

ans is x=0, x=8

OpenStudy (calculator):

I get 2 extra values?

OpenStudy (kropot72):

The coordinates of a stationary point on the curve should be an (x, y) point.

OpenStudy (anonymous):

@Calculator , The mistake you are doing is by squaring the following \[\frac{4}{\sqrt{8x-x^2}}=\frac{x}{\sqrt{8x-x^2}}\] By squaring, you are increasing the number of solutions

OpenStudy (calculator):

how u solve it if dont square

OpenStudy (anonymous):

From your solutions, you can eliminate x = -4 because x(8−x)>=0

OpenStudy (anonymous):

Now why x = 4 should not be your solution is the thing I have to investigate. Hope you understood why x = -4 cannot be your solution

OpenStudy (anonymous):

\[y'=\frac{1}{2}\frac{1}{(8x-x^2)}(8-2x) = 0\] \[y'=\frac{(4-x)}{ \sqrt {(8x-x^2)}} = 0\] So according to me \[x \neq 0,8\] Because if we substitute x=0 or x=8, then y' becomes undeined as denominator becomes zero

OpenStudy (anonymous):

*undefined

OpenStudy (anonymous):

y^2 = 8x-x^2 2y dy = 8-2x dx dy/dx = (4-x)/y zero when x = 4 (so y = 4)

OpenStudy (anonymous):

@estudier , but calculator is saying his solution is x=0,8.. How is it possible. Even I am getting x=4

OpenStudy (anonymous):

See this comment by @calculator "ans is x=0, x=8" "I get 2 extra values?"

OpenStudy (anonymous):

He appears to be treating the equation as an ellipse but equation is only valid for positive x. (The 0 values are minima not maxima)

OpenStudy (anonymous):

Well, I know this is circle equation. But he says the solution given to him is x=0,8 which is puzzling me. According to me x=4 is the only solution to this. Maybe he must be having the wrong solution . Don't know .....

OpenStudy (anonymous):

At 0 and 8 he has minima, not maxima.

OpenStudy (anonymous):

And only if you consider as ellipse which is not valid

OpenStudy (anonymous):

@estudier , you are right that x=0,8 is the point where minima occurs. And maxima occurs at x=4 Now,The stationary point is where the gradient has flatlined,and that point is anywhere the derivative = 0 . So the answer for the question where stationary point is present is x=4, right??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok. :D